OpenStudy (anonymous):

Can any body show me the steps to solve this 8=8sin(165+30a)

OpenStudy (anonymous):

First, you want to move the coefficient in front of the sin function so that sin() is by itself. To do so, you divide by 8. You new equation woudl be 1=sin(165+30a) For reference, sin(f(x)) = 1 \[sin^{-1}(sin(f(x))=f(x)\] \[sin^{-1}(1)=\frac{\pi}{2}\] From there, it's an algebra problem with 1 variable